# Nova: An Iterative Planning and Search Approach to Enhance Novelty and Diversity of LLM Generated Ideas

### v1

@@ Abstract
Xiang Hu4,∗, Hongyu Fu5,∗, Jinge Wang1, Yifeng Wang6, Zhikun Li7 Renjun Xu2, Yu Lu1, Yaochu Jin1, Lili Pan†,3, Zhenzhong Lan†,1 Westlake University1    Zhejiang University2 University of Electronic Science and Technology of China3 China Life R&D Center4    Carnegie Mellon University5    Southeast University6    University of Oxford7 huxiang2022@e-chinalife.comhongyuf@andrew.cmu.edu lilipan@uestc.edu.cnlanzhenzhong@westlake.edu.cn
@@ 2.1 LLM-based Scientific Innovation
In the past year, several studies on LLM-based scientific innovation (Yang et al., 2024; Baek et al., 2024; Lu et al., 2024; Wang et al., 2024b; Gu and Krenn, 2024) have been proposed, garnering significant attention from the LLM community. Among these studies, Baek et al. (2024) introduces a research agent that utilizes an external knowledge graph for co-occurrence entity search and integrates retrieved entities into idea generation of LLMs. To avoid generating similar ideas, Lu et al. (2024) treat past generated ideas as negative examples and instruct the LLM on what constitutes a negative example. To explore more external knowledge for innovation, some other works (Wang et al., 2024b; Gu and Krenn, 2024) propose prompting the LLM to generate ideas integrated with external knowledge, such as retrieved external entities or problem-solution pairs.
@@
Although effective, the above approach often generate repetitive ideas Si et al. (2024) due the lack of direction in acquiring new knowledge. In contract, our method provide a plan for searching new knowledge and suffer less from the repetitive problem.
@@ 3.1 Initial Seed Idea Generation
To produce high-quality ideas, we design a multi-source seed idea generation module that initiates with diverse and novel concepts. This module activates the LLM to generate ideas using related literature and scientific discovery techniques upon receiving an input paper. The prompt for initial idea generation is in Tab. 1 (details is in Tab. 9) and an example of an initial seed idea is in Tab. 2.
@@ 3.2.1 Planning and Search
New Seed Idea Generation. Once new knowledge is acquired, the new seed idea is generated based on the retrieved papers, the initial seed idea, and given input paper. For each idea, our models generate 10 new seed ideas and then use self-reflection to cut the number down to 3.
@@ 3.3 Output Idea Generation
After finishing $T$ step iteration, we have a final seed idea pool. We then expand the seed idea into the initial proposal and final proposal as in Si et al. (2024). Specifically, given an input paper and its corresponding seed idea , we ask LLM to decompose the idea into several sub-modules and utilize LLMs to design these sub-modules separately in a more detailed way(details is in Tab. 32 in Appendix). An example of an initial proposal and final proposal are in Tab. 4 and Tab. 5, separately.
@@ 4.1 Experimental Setup
Data. Our dataset is constructed by collecting high-quality papers from top conferences. The initial corpus comprised 7,805 papers from CVPR 2024, ACL 2024, and ICLR 2024. The keywords related to “LLM" are used to filter the initial corpus down to about 2,000 papers. Minimum citation number is used to further cut down the number to 153, with citation thresholds set at 20 for ICLR 2024 and 10 for CVPR 2024 and ACL 2024. We add an additional 17 papers from Hugging Face Daily Papers according to their user ratings. At the end, the dataset consists of 170 papers, each of which is used to generate 100 ideas for subsequent evaluation.
@@
1. Quality. Following Si et al. (2024), we employ the Swiss System Tournament https://en.wikipedia.org/wiki/Swiss-system_tournament with Claude-3.5-Sonnet zero-shot ranker to evaluate the quality of ideas. The ranker makes pairwise comparisons to determine which idea is better. For each idea, there are 5 rounds of comparison, each winning comparison get 1 score. Such a quality evaluation method has been shown to be better than direct comparison (Lu et al., 2024).
2. Novelty. Following Baek et al. (2024), we use LLMs to judge whether a generated idea is novel by checking the top 10 most relevant papers We use the same paper retriever as in the idea generation phase, with data sourced from arXiv. The time range spans from January 1, 2022, to August 2024, and the categories include literature related to AI, NLP, and CV. and if no paper is identified as containing similar idea, it is considered novel. We use embedding using the all-MiniLM-L6-v2 modelhttps://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 and if the cosine similarity threshold is large than 0.3, we say that they are similar (Si et al., 2024).
@@
In the automatic evaluation, we utilize Nova alongside three baseline methods to generate 400 ideas from a given input paper. Each method generates 100 ideas separately. The iteration step for Nova is set to 3, and the initial number of seed ideas is 15. After 3 iterations, we get 405 initial seed ideas, we then cluster the ideas using k-means clustering into 100 clusters and used the cluster center to generate the 100 final ideas. The implementation details for the baseline methods remain consistent with those in the original work.
Human Evaluation. To validate the effectiveness of our automatic evaluation, we have an additional human evaluation. Our goal is to assess how well our automatic evaluation align with human expert evaluations. We recruit a panel of 10 experts, all holding a PhD degree or professorship in natural language processing, machine learning, or computer vision, doing research in LLMs-related fields. These experts evaluated ideas based on novelty and overall quality (including feasibility and effectiveness).
We select five ideas generated by each agent based on the same input paper. These ideas correspond to the $1^{st}$, $25^{th}$, $50^{th}$, $75^{th}$, and $100^{th}$ percentiles of the automatic evaluation, resulting in a total of $20$ ideas per topic. This process is repeated for $20$ times. Each expert review four groups, ensuring that at least two independent experts evaluate each idea. The final score for each idea is averaged across all ratings from different experts.
We compare the distributions of expert evaluations against those in automatic evaluation. Specifically, we track which methods produce the top 20 percent of ideas, as rank by the experts. This helps determine which methods outperform the others. Moreover, this approach reveals whether the model evaluation aligns with human evaluation.
@@ 4.3 Ablation Study
To assess the effectiveness of planning and search in Nova, we conduct comparisons by gradually removing planning and retrieval components. All methods retrieve the same number of papers, specifically $K=5$. Both retrieval and planning are found to significantly enhance the generation of unique and novel ideas. When planning is excluded, the number of unique ideas at step 3 ($44.1$) is no longer increases compared to step 2 ($42.4$). This suggests that without planning, relying solely on retrieval based on seed ideas limits access to valuable external knowledge for innovation. This limitation may arise from the restricted scope of search when planning is absent. Obviously, when planning and retrieval are both removed, the number of unique novel ideas increases slightly at step 2 (from $25.3$ to $30.6$) and stagnates at step 3 (from $30.6$ to $31.35$), due to no external knowledge are introduced.
@@ 5 Conclusion
In this paper, we propose an LLM-based scientific innovation method, Nova, which introduces iterative planning and search to retrieve external knowledge for innovation. Nova leverages the internal knowledge of LLMs to generate search plans for external knowledge retrieval, significantly enhancing the effectiveness of the retrieval process. Ablation study demonstrates the effect of the iterative planning and search framework on promoting the novelty of generating ideas. The automatic and human evaluations show that Nova significantly and consistently outperforms state-of-the-art scientific innovation methods. In the future, we will explore incorporating a reward function into our iterative planning framework to further enhance external knowledge retrieval.
@@ 6 Limitations
Limited Iterations Steps. Although our approach can significantly enhance the novelty and diversity of generated ideas through iteration, we do not see an continues incrasement in generating new ideas after 3 round of iteration.
@@
We hope these findings inspire future investigations into using LLM to comprehensively integrate both internal and external knowledge for LLM-based scientific innovation. We believe addressing each of these shortcomings will lead to exciting feature directions.
@@ A.1 Annotation Instructions
Annotation Method: Annotate the best 4 and the worst 4 ideas in each of the 4 dimensions. For the best ideas, mark them as 1, 2, 3, and 4 respectively, while 1 refers to the best idea. For the worst ideas, mark them as 17, 18, 19, and 20, while 20 refers to the worsts idea. No need to annotate ideas other than the best 4 and the worst 4. For an example, see Table 7.
@@ A.3 Risk Statement
Social Risk. This study does not involve activities that could affect participants’ social relationships or reputation. No personal information will be disclosed.

### v2

@@ Abstract
Xiang Hu4,∗, Hongyu Fu5,∗, Jinge Wang1, Yifeng Wang6, Zhikun Li7 Renjun Xu2, Yu Lu1, Yaochu Jin1, Lili Pan†,3, Zhenzhong Lan†,1 Westlake University1    Zhejiang University2 University of Electronic Science and Technology of China3 China Life R&D Center4    Carnegie Mellon University5    Southeast University6    University of Oxford7 huxiang2022@e-chinalife.comhongyuf@andrew.cmu.eduwangjinge@westlake.edu.cn lilipan@uestc.edu.cnlanzhenzhong@westlake.edu.cn
@@ 2.1 LLM-based Scientific Innovation
In the past year, several studies on LLM-based scientific innovation (Yang et al., 2024; Baek et al., 2024; Lu et al., 2024; Wang et al., 2024b; Gu and Krenn, 2024; Li et al., 2024) have been proposed, garnering significant attention from the LLM community. Among these studies, Baek et al. (2024) introduces a research agent that utilizes an external knowledge graph for co-occurrence entity search and integrates retrieved entities into idea generation of LLMs. To avoid generating similar ideas, Lu et al. (2024) treat past generated ideas as negative examples and instruct the LLM on what constitutes a negative example. To explore more external knowledge for innovation, some other works (Wang et al., 2024b; Gu and Krenn, 2024) propose prompting the LLM to generate ideas integrated with external knowledge, such as retrieved external entities or problem-solution pairs.
@@
Although effective, the above approach often generates repetitive ideas Si et al. (2024) due to the lack of direction in acquiring new knowledge. In contrast, our method provides a plan for searching for new knowledge and suffers less from the repetitive problem.
@@ 3.1 Initial Seed Idea Generation
To produce high-quality ideas, we design a multi-source seed idea generation module that initiates with diverse and novel concepts. This module activates the LLM to generate ideas using related literature and scientific discovery techniques upon receiving an input paper. The prompt for initial idea generation is in Tab. 1 (details in Tab. 9) and an example of an initial seed idea is in Tab. 2.
@@ 3.2.1 Planning and Search
New Seed Idea Generation. Once new knowledge is acquired, the new seed idea is generated based on the retrieved papers, the initial seed idea, and the given input paper. For each idea, our models generate 10 new seed ideas and then use self-reflection to cut the number down to 3.
@@ 3.3 Output Idea Generation
After finishing $T$ step iteration, we have a final seed idea pool. We then expand the seed idea into the initial proposal and final proposal as in Si et al. (2024). Specifically, given an input paper and its corresponding seed idea, we ask LLM to decompose the idea into several sub-modules and utilize LLMs to design these sub-modules separately in a more detailed way(details in Tab. 32 in Appendix). An example of an initial proposal and final proposal are in Tab. 4 and Tab. 5, separately.
@@ 4.1 Experimental Setup
Data. Our dataset is constructed by collecting high-quality papers from top conferences. The initial corpus comprised 7,805 papers from CVPR 2024, ACL 2024, and ICLR 2024. The keywords related to “LLM" are used to filter the initial corpus down to about 2,000 papers. The minimum citation number is used to further cut down the number to 153, with citation thresholds set at 20 for ICLR 2024 and 10 for CVPR 2024 and ACL 2024. We add additional 17 papers from Hugging Face Daily Papers according to their user ratings. At the end, the dataset consists of 170 papers, each of which is used to generate 100 ideas for subsequent evaluation.
@@
1. Quality. Following Si et al. (2024), we employ the Swiss System Tournament https://en.wikipedia.org/wiki/Swiss-system_tournament with Claude-3.5-Sonnet zero-shot ranker to evaluate the quality of ideas. The ranker makes pairwise comparisons to determine which idea is better. For each idea, there are 5 rounds of comparison, each winning comparison gets 1 score. Such a quality evaluation method has been shown to be better than direct comparison (Lu et al., 2024).
2. Novelty. Following Baek et al. (2024), we use LLMs to judge whether a generated idea is novel by checking the top 10 most relevant papers We use the same paper retriever as in the idea generation phase, with data sourced from arXiv. The time range spans from January 1, 2022, to August 2024, and the categories include literature related to AI, NLP, and CV. and if no paper is identified as containing a similar idea, it is considered novel. We use embedding using the all-MiniLM-L6-v2 modelhttps://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 and if the cosine similarity threshold is larger than 0.3, we say that they are similar (Si et al., 2024).
@@
In the automatic evaluation, we utilize Nova alongside three baseline methods to generate 400 ideas from a given input paper. Each method generates 100 ideas separately. The iteration step for Nova is set to 3, and the initial number of seed ideas is 15. After 3 iterations, we get 405 initial seed ideas, we then cluster the ideas using k-means clustering into 100 clusters and use the cluster center to generate the 100 final ideas. The implementation details for the baseline methods remain consistent with those in the original work.
Human Evaluation. To validate the effectiveness of our automatic evaluation, we have an additional human evaluation. Our goal is to assess how well our automatic evaluation aligns with human expert evaluations. We recruit a panel of 10 experts, all holding a PhD degree or professorship in natural language processing, machine learning, or computer vision, doing research in LLMs-related fields. These experts evaluated ideas based on novelty and overall quality (including feasibility and effectiveness).
We select five ideas generated by each agent based on the same input paper. These ideas correspond to the $1^{st}$, $25^{th}$, $50^{th}$, $75^{th}$, and $100^{th}$ percentiles of the automatic evaluation, resulting in a total of $20$ ideas per topic. This process is repeated for $20$ times. Each expert reviews four groups, ensuring that at least two independent experts evaluate each idea. The final score for each idea is averaged across all ratings from different experts.
We compare the distributions of expert evaluations against those in automatic evaluation. Specifically, we track which methods produce the top 20 percent of ideas, as ranked by the experts. This helps determine which methods outperform the others. Moreover, this approach reveals whether the model evaluation aligns with human evaluation.
@@ 4.3 Ablation Study
To assess the effectiveness of planning and search in Nova, we conduct comparisons by gradually removing planning and retrieval components. All methods retrieve the same number of papers, specifically $K=5$. Both retrieval and planning are found to significantly enhance the generation of unique and novel ideas. When planning is excluded, the number of unique ideas at step 3 ($44.1$) no longer increases compared to step 2 ($42.4$). This suggests that without planning, relying solely on retrieval based on seed ideas limits access to valuable external knowledge for innovation. This limitation may arise from the restricted scope of search when planning is absent. Obviously, when planning and retrieval are both removed, the number of unique novel ideas increases slightly at step 2 (from $25.3$ to $30.6$) and stagnates at step 3 (from $30.6$ to $31.35$), due to no external knowledge being introduced.
@@ 5 Conclusion
In this paper, we propose an LLM-based scientific innovation method, Nova, which introduces iterative planning and search to retrieve external knowledge for innovation. Nova leverages the internal knowledge of LLMs to generate search plans for external knowledge retrieval, significantly enhancing the effectiveness of the retrieval process. The ablation study demonstrates the effect of the iterative planning and search framework on promoting the novelty of generating ideas. The automatic and human evaluations show that Nova significantly and consistently outperforms state-of-the-art scientific innovation methods. In the future, we will explore incorporating a reward function into our iterative planning framework to further enhance external knowledge retrieval.
@@ 6 Limitations
Limited Iterations Steps. Although our approach can significantly enhance the novelty and diversity of generated ideas through iteration, we do not see a continuous increment in generating new ideas after 3 rounds of iteration.
@@
We hope these findings inspire future investigations into using LLM to comprehensively integrate both internal and external knowledge for LLM-based scientific innovation. We believe addressing each of these shortcomings will lead to exciting future directions.
@@ A.1 Annotation Instructions
Annotation Method: Annotate the best 4 and the worst 4 ideas in each of the 4 dimensions. For the best ideas, mark them as 1, 2, 3, and 4 respectively, while 1 refers to the best idea. For the worst ideas, mark them as 17, 18, 19, and 20, while 20 refers to the worst idea. No need to annotate ideas other than the best 4 and the worst 4. For an example, see Table 7.
@@ A.3 Risk Statement
Social Risk. This study does not involve activities that could affect participants’ social relationships or reputations. No personal information will be disclosed.